OpenStudy (anonymous):

The world’s tallest unsupported flagpole is a 282-ft-tall steel pole in Surrey, British Columbia. The shortest shadow cast by the pole during the year is 137 ft long. To the nearest degree, what is the angle of elevation of the sun when casting the flagpole’s shortest shadow? 1. Draw a sketch using the information given in the problem. Draw an arrow pointing to the angle of elevation.

OpenStudy (anonymous):

Can someone help?

OpenStudy (anonymous):

@Palmo4ka ?

OpenStudy (anonymous):

@lovatic4life

OpenStudy (anonymous):

 \[tg \alpha= 282/137 \approx 2\] the angle = 244 degrees

OpenStudy (anonymous):

Thanks, I have more questions if you dont mind @Palmo4ka

OpenStudy (anonymous):

sorry, i have to go now, but i will not mind to a medal :D

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

@stassiexoxo ?

OpenStudy (anonymous):

@beccaboo333 help?

OpenStudy (anonymous):

@linn99123 ?